OpenStudy (anonymous):

60/56 as a mixed number

jagr2713 (jagr2713):

Reduce the expression 60/56 by removing a factor of 4 from the numerator and denominator.

jagr2713 (jagr2713):

so its 15/14

OpenStudy (anonymous):

1 and 4/56

OpenStudy (anonymous):

\[1+\frac{1}{14} \]

OpenStudy (anonymous):

60= 1*56 +4

OpenStudy (anonymous):

First, divide the numerator (the 60) by the denominator (the 56). This will give you 1, with a remainder of 4.

hero (hero):

\[\frac{60}{56} = \frac{56 + 4}{56} = \frac{56}{56} + \frac{4}{56} = 1 + \frac{1}{14}\]

OpenStudy (anonymous):

so it will be q/d so q=a*d +r 60/56 60=1*56 +4